40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 21 19:28:03 2021, Anchal, Update, PSL, PSL/IFO recovery 
    Reply  Wed Apr 21 19:43:20 2021, Koji, Update, PSL, New HEPA speed control P_20210421_193637.jpgP_20210421_193627.jpg
       Reply  Thu Apr 22 14:41:55 2021, Chub, Update, PSL, New HEPA speed control 40M_PSL_HEPA.jpg
          Reply  Fri Apr 23 18:00:02 2021, gautam, Update, PSL, HEPA speed lowered HEPAdiag.pdf
             Reply  Fri Apr 23 19:26:58 2021, Koji, Update, PSL, HEPA speed lowered 
             Reply  Fri May 14 17:45:05 2021, rana, Update, PSL, HEPA speed raised 
                Reply  Tue May 18 00:52:38 2021, rana, Update, PSL, HEPA speed raised Untitled.png
                   Reply  Tue May 18 20:26:11 2021, rana, Update, PSL, HEPA speed raised 
    Reply  Thu Apr 22 15:15:26 2021, gautam, Update, PSL, PMC transmission 
       Reply  Thu Apr 22 15:34:54 2021, Anchal, Update, PSL, PMC transmission 
Message ID: 16068     Entry time: Wed Apr 21 19:28:03 2021     Reply to this: 16069   16076
Author: Anchal 
Type: Update 
Category: PSL 
Subject: PSL/IFO recovery 

[Anchal, Koji]

Removed the top sheet

  • Opened first from the door side so that any dust would spill outside.
  • Then rolled the sheet inward to meet in the middle.
  • Repeated this twice for the 2 HEPA filters.

Removed the sheets on the table

  • Lifted sheet up making sure the top side face outside always.
  • Rolled it sideways halfway through.
  • Cut down the sheet vertically.
  • Slided the doors to the other side and rolled the remaining half.
  • On the door side, the sheets above the ALS optics were simply lifted off.

Restarting PSL

  • Turned on the HEPAs at the max speed
  • Switched on laser to jsut above the threshold
    • Before the 1st eom, power was 20mW 
    • After the EOM/AOM, 18mW. So about 90% transmission through all polarizing optics.
    • We saw the resonances of the PMC but could not lock it even with highest gain available (30 dB).
  • Increased the input power to PMC to 100mW
    • Locked the PMC at 30dB gain
    • The transmitted power was ~50-60 mW. (Had to use power meter suspended by hand only.
    • The right before the IMC (after the 2nd EOM) 48mW. So none of the alignment was lost.
  • Opened the PSL shutter.
  • We were able to see IMC reflection signal.
  • We were also able to see IMC catching lock as the servo was left ON earlier.
  • Switched off the servo.
  • Decided to increase the power while watching PMC Trans/Refl and IMC REFL
  • Injection diode current to innolight was increased slowly to 2.10A. Saw a mod hopping region aroun 1.8A.
  • We recovered the PMC Trans >0.7 V.
  • PZT was near the edge, so moved by one FSR.
  • The PMC refelction signal is still shown in red at around 48 mV.

Back to control room

  • IMC was locked almost immediately by manually finding the lock while keeping IMC WFS off to preserve the offsets from yesterday.
  • Then switch on IMC WFS. Working good.
  • Then unlocked the servo and switched on IMC Autolocker. Lock was caught immediately.

Decided to start locking the arms

  • The arm transmissions were flashing but at 0.2~0.3 level.
  • Decided to adjust TT1 and TT2 Pitch and Yaw to align the light going into the arms.
  • This made TRY ~0.6 / TRX ~0.8 at the peak of the flashing
  • Locked the arms. (By switching on C1:LSC-MODE_SELECT which engages all servos).
  • Used ASS to align Yarm then align Xarm. Procedure:
    • Sitemap > ASC > c1ass
    • Open striptool to look at progress. ! Scripts YARM > striptool.
    • Switch on ASS. ! More Scripts > ON
    • Wait for the TRY to reach to around 0.97.
    • Freeze the outputs. ! Scripts > Freeze Outputs.
    • Offload the offsets to preserve the output. ! More Scripts > OFFLOAD OFFSETS.
    • Switch off ASS. ! More Scripts > OFF
    • Repeted this for XARM.
  • At the end, both XARM and YARM were locked with TRX ~ 0.97 and TRY ~ 0.96.
ELOG V3.1.3-